OpenStudy (anonymous):

Mean Value Theorem: Let f(x)=(x-3)^(-2). Show that there is no value of c in (1,4) such that f(4)-f(1)=f'(c)(4-1). Why does this not contradict the Mean Value Theorem?

OpenStudy (anonymous):

Where is \(f(x)\) continuous?

OpenStudy (anonymous):

I think that's my problem - I don't know how to tell whether it's continuous or not unless it's graphed

OpenStudy (anonymous):

\[f(x)=(x-3)^{-2}=\frac{1}{(x-3)^2}\] What would happen is \(x=3\) ?

OpenStudy (anonymous):

Oh! It wouldn't exist

OpenStudy (anonymous):

So rule of thumb is that if x is in a position for denominator to be zero then it is not continuous

OpenStudy (anonymous):

Yes, that's a way of thinking of it. A vertical asymptote occurs at \(x=3\), if you were to look at the function's plot. So why can't you apply the MVT?

OpenStudy (anonymous):

Because 3 is within (1,4) but it DNE

OpenStudy (anonymous):

Right. More specifically, a function has to be continuous over a given interval. This \(f(x)\) is not, and so MVT does not apply.

OpenStudy (anonymous):

Perfect, thanks for your help!

OpenStudy (anonymous):

you're welcome